OpenStudy (anonymous):

Medal&Fan Alfred and Ashley are changing their rectangular storage closet into a game room. They decide to add 2x feet to the length and width of the storage closet. Which equation will show the new area of the storage closet if it was initially 8 feet by 5 feet? 4x2 + 13x + 13 square feet 4x2 + 13 square feet 4x2 + 40 square feet 4x2 + 26x + 40 square feet

OpenStudy (hoblos):

if the length was 8 feet, and we add 2x feet what would the new length be ?

OpenStudy (anonymous):

16x^2

OpenStudy (anonymous):

?idk i think

OpenStudy (hoblos):

8 add 2x to it, it will be 8+2x what about the width ?

OpenStudy (anonymous):

5+2x

OpenStudy (anonymous):

soo its C?

OpenStudy (hoblos):

the area of a rectangle is lenght times width area = (8+2x)(5+2x) can you expand this ?

OpenStudy (anonymous):

8*5=40 2X*2X = 4X^2 so 4x^2+40

OpenStudy (anonymous):

shoot C isnt right

OpenStudy (anonymous):

its D

OpenStudy (hoblos):

(8+2x)(5+2x) = 8(5) + 8(2x) + 2x(5) + 2x(2x) = 40 + 16x + 10x + 4x2 = 4x2 + 26x +40 so yeah it's D

OpenStudy (anonymous):

Thanks man.(: